Find the equation to the straight line passsing through the point (3,-4) and cutting of intercepts equal but of opposite signs from the two axis

Grade:10

1 Answers

Aditya Gupta
4 years ago
eqn will be y+4= m(x – 3)
now, intercepts are (3+4/m, 0) and (0, –(3m+4))
they are equal and opposite.
so – (3+4/m)= –(3m+4)
or 4/m= 3m + 1
so, m= 1, – 4/3
but when m= – 4/3, both intercepts are 0. 
so, m=1
eqn becomes
y+7= x
kindly approve :)